# Trajectory Prediction for Autonomous Driving: Progress, Limitations, and Future Directions

### v1

@@ 9.1 On Research Question 1
Answering question 1, trajectory prediction has evolved through several distinct phases of modelling approaches, driven by advances in deep learning and machine learning techniques. Early deep-learning methods were heavily reliant on recurrent models, such as RNNs and LSTMs, which effectively captured temporal dependencies in sequential data. Around 2017-2018, these models gained significant attention for their ability to process dynamic motion patterns, often through hierarchical or multi-layer designs. As the field progressed, modeling more complex agent interactions led to the emergence of attention-based mechanisms and graph-based approaches.
The current trend shifts towards Transformer architectures and Graph Neural Networks (GNNs), which have emerged as SOTA since 2020. Transformers leverage self-attention to model both spatial and temporal dependencies across agents, offering adaptability in heterogeneous traffic environments. GNNs, on the other hand, excel in capturing agent-to-agent interactions through graph representations, enhancing their effectiveness in complex, multimodal scenarios.
SOTA models commonly integrate various architectures to address the complexities of trajectory prediction. These include CNNs for scene understanding, attention mechanisms for feature weighting, and encoder-decoder setups for trajectory generation. Examples of such combinations include graph-enhanced Transformers and CNN-LSTM pipelines. Additionally, hybrid solutions that combine physics-based modelling with deep learning approaches are gaining attention for their ability to integrate domain knowledge with data-driven adaptability; like, models incorporating kinematic constraints alongside neural networks enhance prediction accuracy while maintaining physical plausibility.
@@ 9.2 On Research Question 2
Answering question 2, recent research trends in the prediction domain focus on enhancing the robustness and generalization of prediction models. As discussed in Section 5, a significant research gap exists in quantifying and mitigating uncertainty within prediction pipelines. One approach to addressing aleatoric uncertainty is through collaborative perception, which compensates for individual perception system limitations, such as restricted sensing capabilities, occlusions, and sensor noise. Collaborative perception enables multiple vehicles to share information, creating a more comprehensive environmental view. This area, detailed in Section 8, is investigated in detection and tracking and is gradually expanding into prediction, underscoring the need for further investigation and the development of collaborative prediction models.
In improving robustness and generalization, formalizing and incorporating driving knowledge remains an active area of research, as outlined in Section 4. Planning-conditioned trajectory prediction methods, as discussed in Section 6, represent a promising direction by incorporating motion planning principles into prediction frameworks. These methods assume rational agent behavior and condition predictions on agents’ motion plans or intended goals. By modelling planned trajectories and potential destinations, planning-conditioned approaches provide interpretable and constraint-compliant predictions, particularly in interactive and complex scenarios. Additionally, Vision-Language Models (VLMs), as discussed in Section 7, offer common knowledge and contextual understanding necessary for navigating complex and uncertain environments. While VLMs hold promise as anchors in uncertain conditions, their real-world application in autonomous driving is still under evaluation. Practical challenges, such as hosting these large models on the cloud, can introduce latency issues that may compromise real-time predictions. The integration of VLMs into the autonomous driving stack requires further exploration to determine their effectiveness and deployment configurations.
These trends reflect a shift in trajectory prediction research towards developing robust, collaborative, context-aware, and planning-informed methods that prioritize both accuracy and real-time operation in dynamic environments.
@@ 9.3 On Research Question 3
On question 3, in the overview of the prediction pipeline, Sections 2.6 and 2.7 outline the metrics used for evaluating prediction models and list benchmark datasets with leaderboard solutions. A challenge in comparing different methods from the literature (apart from leaderboard rankings) lies in inconsistencies in the prediction horizon across the same dataset. Additionally, comparing methods based on performance across multiple datasets is difficult, as most methods are evaluated on only one or, at most, two datasets. Few studies address the critical issue of robustness, which is essential to ensure reliable predictions despite detection errors, tracking deficiencies, localization inaccuracies, or map changes. Furthermore, no research has systematically compared methods under such real-world conditions.
Despite the extensive range of methods and refined solutions, it is evident that the rate of improvement in prediction accuracy is diminishing. On leaderboards, the performance gaps between top solutions show only marginal gains. In practical scenarios, extremely precise trajectory forecasts may not significantly enhance the effectiveness of prediction modules. Human drivers, while less precise than autonomous vehicles, navigate effectively across varying traffic densities. Thus, understanding how such high precision contributes to the overall success of an autonomous stack, including collision avoidance, is crucial. While prediction models achieve higher precision in most scenarios, failures in specific situations could lead to severe consequences. Current studies lack qualitative evaluations to identify scenarios where predictions falter, and there is a noticeable gap in testing for generalization. Specifically, cross-dataset evaluations, where models trained on one dataset are tested on another, are underexplored, leaving performance in out-of-distribution scenarios unaddressed. Consequently, while existing models demonstrate strong performance using reported metrics, these evaluations often fail to comprehensively reflect real-world operational needs, requiring further testing and investigation.
@@ 9.4 Future Directions
9.4.1 Real-time Evaluation
For autonomous vehicles to navigate complex traffic environments safely, they require a deep semantic understanding of both static and dynamic scene elements. Achieving this understanding necessitates advanced deep learning architectures with multiple processing and reasoning layers to interpret contextual features, predict agent behavior, and plan appropriate responses. However, the incorporation of semantic layers to enhance comprehension inevitably increases inference time, presenting a critical trade-off between model complexity and processing speed.
Lin et al. [642] emphasize that real-time autonomous vehicle systems must process and react to traffic conditions faster than human drivers to ensure safety. While the fastest human reaction times range from 100–150ms [643, 644], autonomous systems are required to achieve processing latencies below 100ms and maintain a frame rate exceeding 10 frames per second. This dual requirement enables both rapid responses to individual events and continuous awareness of dynamically changing traffic conditions. These stringent latency demands have significant implications for the design of autonomous driving pipelines. In a complete pipeline encompassing detection, tracking, prediction, and planning, the combined latency must remain under 100ms, establishing a strict upper bound for the prediction module. Despite this, much of the current research focuses on maximizing prediction accuracy while neglecting real-world inference-time evaluation, creating a gap between theoretical advancements and practical deployment. Real-world applications demand a balance between accuracy and speed to enable seamless integration of prediction systems with planning modules, underscoring a critical area for further development.
When designing prediction systems, it is essential to prioritize architectures that can be effectively parallelized and fully leverage available hardware to minimize inference time. Attention-based models, in particular, offer significant advantages in this regard. Their parallel computation capabilities enable faster inference times on modern GPUs and TPUs optimized for such operations. While factors such as model size, sequence length, and memory requirements affect absolute performance, the parallel nature of attention mechanisms allows for further optimization through techniques like multi-head parallelization and batch processing. This advantage has been demonstrated in studies such as [645, 245], where optimized embedding systems combined with hardware-aware implementations significantly improved inference speed without sacrificing prediction accuracy.
Model selection should be guided by the hardware capabilities of the deployment platform and the system’s operational requirements. In some scenarios, a lighter, faster model (e.g., attention-based architectures with optimized implementations) that provides near-real-time predictions may be more practical than a more accurate but slower alternative. While such models may not achieve state-of-the-art accuracy on benchmarks, their ability to deliver rapid updates enables more frequent error correction and quicker system responses. In dynamic traffic scenarios, where predictions must be continuously refined, the capability to quickly adapt to changes is often more critical for safe navigation than marginally higher accuracy achieved at the cost of computational speed.
9.4.2 Enhancing Robustness
Future research must address key challenges to ensure the effective deployment of trajectory prediction systems in real-world autonomous driving environments. Most existing methods are tailored for specific tasks, scenarios, or motion types, performing well under controlled conditions with clear motion patterns, well-defined spatial structures, and predictable objectives. However, these models often falter in dynamic and unpredictable scenarios, highlighting the need for comprehensive evaluation across diverse and realistic conditions.
To achieve robustness, prediction systems must dynamically adapt to environmental changes and variations in agent behavior. Current models are predominantly trained on datasets from limited driving scenarios, which fail to capture the breadth of real-world diversity in driving styles and configurations. Enhancing adaptability to novel and unseen situations remains a critical challenge. Incorporating driving knowledge, such as traffic rules, road layouts, and interaction dynamics, can provide models with a strong foundation for understanding traffic systems, thereby improving their robustness and adaptability. However, managing rare and unexpected behaviors requires a different approach. Drawing inspiration from human cognition, autonomous vehicles (AVs) could adopt a dual-process strategy: relying on predefined, optimized algorithms in familiar scenarios while leveraging generalized reasoning mechanisms for unfamiliar or uncertain situations. Vision-language models (VLMs) show promise in enabling AVs to handle complex, unstructured edge-cases, and research in this area is rapidly advancing, warranting further exploration.
Another critical area gaining attention is the mitigation and quantification of uncertainties, specifically aleatoric (inherent randomness) and epistemic (model-related) uncertainties. These uncertainties must be explicitly addressed and conveyed to downstream planning systems, allowing them to make informed decisions. Despite progress in some models, effectively mitigating uncertainty remains a significant gap in current prediction pipelines. Many existing methodologies assume error-free detection and tracking, often overlooking the impact of perception errors on performance. Evaluation metrics like average displacement error (ADE) are commonly used, reflecting the accuracy of prediction model, but is not sufficient to comprehensively capture the nuances of failure cases or scenarios influenced by perception errors. This narrow focus neglects qualitative insights into critical edge cases where prediction failures could result in severe consequences, such as collisions. Future evaluation frameworks should emphasize a balance between robustness and accuracy. While minor displacement errors can be mitigated through conservative planning strategies, complete prediction failures pose a substantial safety risk. Robustness under perception errors must be prioritized, with evaluation strategies focusing on failure case analysis to build more resilient and reliable systems capable of navigating real-world complexities.
Additionally, most current models specialize in either short- or long-horizon predictions, neglecting the benefits of integrating both temporal scales. Long-horizon predictions are essential for strategic planning, enabling proactive trajectory optimization and efficient navigation. However, their inherent uncertainty increases with the prediction horizon. Conversely, short-horizon predictions offer precise immediate forecasts for rapid reactive measures but lack the broader context required for long-term planning. A more effective solution involves developing integrated prediction systems that combine the strengths of both temporal scales. In such systems, long-horizon predictions establish a strategic plan, continuously refined by short-horizon predictions in real-time. This iterative refinement creates a robust and adaptive framework capable of both proactive decision-making and reactive safety responses, addressing the complexities of real-world driving while ensuring safety and efficiency.
@@ 9.4.2 Enhancing Robustness
Finally, expanding a model’s capacity to handle novel scenarios beyond its initial training distribution is essential for safety. This issue can be addressed through simulation-to-reality (Sim2Real) adaptation and continual learning. Sim2Real techniques involve generating synthetic scenes mimicking real-world scenarios and conditions, allowing models to learn from controllable environments, where different levels of diversity can be gradually introduced before deploying the model in actual settings. For continual learning, during trial deployments or pre-deployment phases, a small dataset specific to the target city can be collected to adapt the model to local conditions. This continual learning phase ensures the model is fine-tuned to the deployment context while retaining prior knowledge. Although some works, such as [646, 647, 648], explore this direction, further research is needed to fully realize its potential in autonomous systems.
9.4.3 Heterogeneous Trajectory Prediction
In autonomous driving, trajectory prediction must account for all traffic agents, including vehicles of varying sizes (e.g., trucks, vans, sports cars), pedestrians, bicycles, and motorcycles. Each of these agents exhibits unique kinematics, significantly influencing their motion patterns. For vehicles, kinematic constraints, such as steering limits and acceleration bounds, heavily restrict possible trajectories. In contrast, pedestrian motion is far less constrained, introducing greater randomness and variability. This distinction makes pedestrian trajectory forecasting particularly challenging, as it is influenced by a broader range of factors, including individual intent, environmental context, and social interactions. A critical challenge lies in designing a unified prediction framework that accommodates the diverse characteristics of all agent types. One approach is to use a single model trained on extensive data to learn the dynamics of each agent class implicitly. While this enables scalability and uniformity, it often limits the model’s ability to explicitly integrate kinematic constraints, especially for agents with well-defined physical dynamics, such as vehicles. On the other hand, a specialized approach involves using separate prediction models tailored to specific agent classes, enabling the explicit incorporation of agent-specific constraints and behaviors. However, this method introduces complexity, including the need to accurately classify agents and manage potential detection and classification errors. In both approaches, the reliance on object classification and detection performance introduces dependencies that must be carefully addressed. Misclassification or detection errors can propagate through the system, degrading prediction accuracy. Therefore, robust mechanisms for error handling and uncertainty modelling are essential. Designing a comprehensive prediction framework that effectively integrates driving knowledge, accounts for the diverse nature of traffic agents, and mitigates detection-related challenges remains an open area of research. Such a framework must strike a balance between generalization and specialization, ensuring safety, accuracy, and scalability in real-world autonomous driving scenarios.
An additional challenge in human trajectory prediction lies in balancing individual and group-aware predictions. For individual motion, it is crucial to incorporate appearance-based features, posture detection, individual characteristics, and elements of behavioral analysis to enhance prediction accuracy. In contrast, group-aware predictions require modelling group formations, relationships, and dynamics to understand collective motion patterns effectively. Previous studies have highlighted the importance of these elements in understanding pedestrian group behavior and dynamics [649, 650, 651, 652, 653, 654, 655, 656, 657, 658]. Developing hybrid models that integrate crowd modelling with individual predictions represents a promising direction, especially in urban environments with dense pedestrian crossings. If a model can accurately recognize crowd formations and adapt by predicting the collective flow and motion of the crowd, computational resources can be significantly conserved. In such scenarios, the detector would not need to identify each individual or run predictions for every agent separately. Conversely, for people who leave the crowd or are not part of it, the system should seamlessly switch to individual-level predictions. This approach requires further investigation, particularly in the development of diverse and representative crowd datasets for autonomous driving applications. Existing datasets [159], serve as a foundation but must be expanded to include more complex and dynamic scenarios. Advancing this area of research will enhance the ability of autonomous systems to operate efficiently and safely in environments characterized by dense and dynamic pedestrian activity.
9.4.4 Collaborative Perception
Collaborative perception enhances the sensing capabilities of individual vehicles by integrating data from multiple sources and viewpoints. This integration allows for more accurate detection of distant or obscured objects and supports robust decision-making in dynamic environments. However, several critical challenges remain to be addressed. One key research direction is ensuring the feasibility of collaborative models in terms of inference time and computational resources. Another crucial aspect is the generalization of collaborative models. Learning-driven approaches excel at deriving comprehensive representations from complementary views, but their ability to generalize to new datasets and scenarios requires thorough investigation and evaluation.
Different collaboration modes introduce unique challenges that require tailored solutions. Vehicle-to-Vehicle (V2V) communication is crucial for enabling rapid and dynamic collaboration by facilitating the sharing of localized sensor data among vehicles. Future research should focus on optimizing data-sharing efficiency under bandwidth constraints and high-speed conditions. Developing context-aware protocols that dynamically determine which information to aggregate and transmit based on real-time conditions is a promising direction. Vehicle-to-Infrastructure (V2I) communication adds a centralized component to collaborative perception by integrating data from fixed infrastructure sensors, such as roadside units and traffic cameras. These sensors offer a broader sensing range and a more holistic perspective on the driving environment. Adapting V2V algorithms for V2I scenarios could improve performance, especially in diverse and complex urban settings. Additionally, research should explore the scalability and resilience of V2I systems, ensuring their reliability despite sensor failures or communication interruptions. Vehicle-to-Everything (V2X) communication, which combines V2V and V2I systems, introduces added complexity due to challenges such as ensuring data authenticity, establishing trustworthiness, and synchronizing information across multiple, diverse entities. To validate the feasibility of V2X systems, future research must extend beyond simulated environments and emphasize testing with real-world data. Experimentation in live traffic scenarios is essential to address these challenges, ensuring the reliability, scalability, and practical applicability of V2X communication in real-world traffic systems.
Finally, designing a robust testing platform is crucial for enabling fast prototyping and evaluation of cooperative driving algorithms. One notable solution is the open-source OpenCDA ecosystem, introduced by Xu et al. [659]. This platform integrates a range of SOTA models for collaborative detection, offering comprehensive support for testing in various cooperative scenarios. While OpenCDA provides tools for benchmarking cooperative perception and driving automation, its extensibility can further be enhanced to address emerging challenges. These include testing benchmarks for communication robustness, V2I collaboration, and diverse real-world scenarios to ensure broader applicability in autonomous driving research.

### v2

@@ 9.1 On Research Question 1
What are the existing solutions for trajectory prediction?
Section 3 showed that trajectory prediction realm has evolved through several distinct modeling phases, starting from physics-based methods and advancing to deep learning hybrid architectures. Existing solutions span classical machine learning, recurrent networks, generative models, graph-based architectures, attention mechanisms, and hybrid methods. Early deep learning approaches relied on recurrent models to capture temporal dependencies in sequential data. The next phase focused more on dynamic motion patterns, often using hierarchical or multi-layer structures. As the field matured, modeling complex patterns and agent interactions led to the emergence of attention mechanisms and graph-based methods.
Current SOTA models commonly integrate multiple architectures to handle the inherent complexity of trajectory prediction, e.g., using CNNs for scene understanding, attention for feature weighting, and encoder–decoder structures for trajectory generation. Examples of such combinations include graph-enhanced Transformers and CNN–LSTM pipelines. In parallel, hybrid approaches that fuse physics-based modeling with deep learning are increasingly explored, integrating domain knowledge with data-driven flexibility. Models that embed kinematic constraints within neural networks enhance prediction accuracy while preserving physical plausibility.
@@ 9.2 On Research Question 2
Is the trajectory prediction problem solved, or is there still room for improvement?
Despite the extensive range of methods and increasingly refined solutions, the rate of improvement in prediction accuracy appears to be diminishing. As shown in Section 2.7, benchmark leaderboards for some SOTA datasets reveal only marginal performance gains between top-performing models. For example, on the Waymo Motion Forecasting dataset, the difference in minADE between the top two solutions is just 0.01 (MTR_v3: 0.55 vs. ModeSeq: 0.56). In real-world applications, extreme precision at the millimeter level does not necessarily guarantee the effectiveness of prediction modules in all scenarios. Human drivers, though less precise, navigate complex environments successfully. While current models demonstrate high averaged accuracy on dataset scenarios, failures in specific edge cases may lead to severe consequences.
Thus, from a deployment readiness perspective, several gaps remain. Most studies lack qualitative evaluations that reveal where models fail, since metrics are typically averaged, making it difficult to extract insights about edge cases. If prediction instability occurs at certain timesteps, introducing a consistency mechanism becomes a relevant research direction. For underrepresented patterns and scenarios, more studies on cross-dataset evaluations are needed, i.e., training on one dataset and testing on another, to assess performance in out-of-distribution scenarios, which remains largely unexplored. Additionally, robustness to upstream errors (e.g., detection, tracking, localization, map quality) is insufficiently addressed in the academic community, despite its critical role in ensuring reliable real-world performance.
@@ 9.3 On Research Question 3
What are the emerging research trends in trajectory forecasting realm?
Recent research trends in the prediction domain focus on enhancing the robustness and generalization of prediction models. As discussed in Section 5, a significant research gap exists in quantifying and mitigating uncertainty within prediction pipelines. One approach to mitigate aleatoric uncertainty is through collaborative perception, presented in Section 8, which compensates for individual perception system limitations through collaborative holistic view of the environment. Planning-conditioned trajectory prediction methods, presented in Section 6, represent a promising direction by incorporating motion planning principles into prediction frameworks. By modeling planned trajectories and potential destinations, planning-conditioned approaches provide interpretable and constraint-compliant predictions, particularly in interactive and complex scenarios.
Additionally, Vision-Language Models (VLMs), as discussed in Section 7, offer common knowledge and contextual understanding necessary for navigating complex and uncertain environments. While utilizing VLMs is a highly emerging and active research area, and holds promise as anchors in uncertain conditions, their real-world application in autonomous driving in terms of latency and predictability is still under investigation. Practical challenges, such as hosting these large models on the cloud, can introduce latency issues that may compromise real-time predictions. The integration of VLMs into the autonomous driving stack requires further exploration to determine their effectiveness and deployment configurations.
These trends reflect a shift in trajectory prediction research towards developing robust, context-aware, and planning-informed methods for operation in dynamic environments.
@@ 9.4 Future Directions
  1. 1.
    Real-time Inference.
    Accurate trajectory prediction in complex traffic scenarios requires semantic understanding of the environment and interpretation of agent behavior. Incorporating advanced reasoning layers to achieve this increases inference time. Lin et al. [649] emphasize that autonomous vehicle systems must process and react to traffic conditions faster than human drivers to ensure safety. While the fastest human reaction times range from 100–150 ms [650, 651], the safe response time for autonomous systems should not exceed 100 ms. This sets the upper limit and defines the latency constraints for the entire perception pipeline, including prediction systems. Yet, much of the current research prioritizes accuracy and lacks inference-time analysis, creating a gap between theoretical advances and practical deployment. Real-world integration calls for design of predictors that balance precision and speed, supplied with inference-time testing.
  2. 2.
    Hardware Compatibility.
    Prediction model selection should be guided by the hardware constraints of the deployment platform and the operational requirements of the system. In many scenarios, a lighter and faster model, like architectures that can be parallelized or hardware-accelerated, may offer lower latency and be more practical than a slower, higher-accuracy alternative. For example, parallelizable models with optimized embeddings, as shown in Giuliari et al. [652], Mebrahtu et al. [242], have demonstrated improved inference speed without compromising performance. While such models may rank lower on benchmark leaderboards, their ability to provide rapid updates allows for more frequent error correction and faster system responses. However, current research primarily focuses on maximizing accuracy, often overlooking evaluation on edge devices or analysis of deployment feasibility under varying resource constraints. This gap hinders the transition from high-performance prototypes to real deployable systems.
  3. 3.
    Integration with Perception.
    As noted in subsection 2.2, in the detect-track-predict paradigm, the input to the prediction pipeline consists of tracklets that inherently carry detection and tracking errors. However, most existing methods assume error-free past trajectories during inference, which is an unrealistic assumption. A key research gap lies in evaluating the robustness of prediction models under perception noise, and in designing tolerance mechanisms or training strategies that allow models trained on clean data to generalize to noisy or incomplete inputs.
  4. 4.
    Prediction for Multiple Horizons.
    Most trajectory prediction models focus on single horizon predictions. Short-term predictions (typically up to 1, 1.5 or 2 seconds) — are essential for handling abrupt events such as sudden braking, sharp lane changes, or obstacle emergence. Errors in this window can escalate quickly, so precise, low-latency forecasts are critical for accounting for random behaviors and maintaining vehicle safety and stability. In contrast, long-horizon predictions (>> 2 seconds) guide strategic decisions like merging or route planning but suffer from growing uncertainty over time. Recent approaches like DEMO [653] adopt a multi-horizon structure that fuses dynamics-informed short-term forecasting with interaction-aware long-term reasoning. These multi-scale predictions can be jointly consumed by planning modules for real-time refinement.
  5. 5.
    Mitigating Epistemic Uncertainty.
@@
    Handling out-of-distribution scenarios remains an active research area. One promising direction involves a dual-process strategy: using predefined, optimized algorithms in familiar scenarios while leveraging the generalized reasoning capabilities of Vision-Language Models (VLMs) in unfamiliar or uncertain situations. Another approach focuses on simulation-to-reality (Sim2Real) adaptation. Sim2Real techniques involve generating synthetic scenes that mimic real-world conditions, enabling models to learn in controllable environments where varying levels of diversity can be introduced before deployment. Additionally, continual learning during trial deployments or pre-deployment phases can help tailor models to local conditions. A small dataset specific to the target city may be collected to fine-tune the model while preserving previously learned knowledge. Although some works, such as [654, 655, 656], explore this direction, further research is needed to fully realize its potential in autonomous systems.
  6. 6.
    Heterogeneous Trajectory Prediction.
    Autonomous vehicles must predict trajectories for diverse agent categories, including vehicles of various sizes (e.g., trucks, vans, sports cars), pedestrians, bicycles, and motorcycles. Each agent type exhibits distinct kinematic properties, typical speeds, and acceleration patterns, which significantly influence their motion behavior. Vehicle trajectories are constrained by factors such as steering limits and acceleration bounds, whereas pedestrian motion is more unconstrained and inherently stochastic. A key challenge is developing a unified prediction framework capable of accommodating these heterogeneous characteristics. One approach involves using a single model trained on large, diverse datasets, with agent category labels and relevant priors provided as inputs, or following a hierarchical design where the model first classifies the agent type and then delegates prediction to category-specific modules. This setup requires carefully curated training data to ensure balanced learning across categories. Alternatively, using category-specific predictors can improve specialization but increases inference time and system complexity.
  7. 7.
    Pedestrian Individual and Group-aware Trajectory Prediction.
    Group-aware predictions require modeling group formations, relationships, and dynamics to effectively capture collective pedestrian motion patterns. Prior studies have emphasized the significance of these factors in understanding group behavior [657, 658, 659, 660, 661, 662, 663, 664, 665, 666]. A promising direction lies in developing hybrid models that integrate crowd-level modeling with individual trajectory prediction, especially in dense urban scenes. If the model can accurately detect group formations and estimate collective motion, it may reduce computational costs by avoiding per-agent predictions within the crowd. For pedestrians outside group formations or leaving the crowd, the system should switch to individual-level predictions. This approach requires further research, particularly in building diverse and representative crowd datasets tailored to autonomous driving contexts.
  8. 8.
    Integration with Planning.
    The output of the prediction module is passed to the planning block in the autonomous driving stack. However, these two components are often developed by separate communities, leaving the integration aspect [667] relatively underexplored. Key integration challenges include suitable predictor’s output format (e.g., occupancy grid is least challenging for integration), quantifying uncertainty, and determining the level of prediction error the planner can tolerate. Ultimately, the goal of prediction is to support safe planning through collision avoidance, thus, tighter integration with planning modules can provide valuable insight into performance requirements and system robustness. To enable smoother interaction, planners must be aware of the uncertainty in predicted trajectories. One effective approach is to use multimodal predictions, which present several possible future outcomes along with their associated probabilities. This allows the planner to consider alternative scenarios and make more informed, flexible decisions when selecting a safe course of action.
  9. 9.
    Collaboration between AVs.
    To overcome limitations in individual sensor suites, collaboration between vehicles enables the construction of more comprehensive scene representations from complementary viewpoints. One promising research direction is task-specific collaboration, where context-aware fusion is performed using sparse, semantically relevant features tailored to the requirements of downstream tasks. Another research direction to explore is the assumption that shared information is always reliable. However, a vehicle with degraded or malfunctioning sensors may unintentionally transmit erroneous data. Therefore, fusion mechanisms should incorporate verification steps to evaluate the credibility of incoming information and avoid propagating low-quality inputs.
  10. 10.
    V2V, V2I and V2X systems.
    Different collaboration modes pose distinct challenges requiring tailored solutions. V2V communication supports rapid, localized data sharing but demands future research on bandwidth-efficient, context-aware protocols that prioritize and transmit only task-relevant information in real time. V2I communication introduces centralized sensing via infrastructure (e.g., roadside units), offering a broader view of the environment. Research should focus on adapting V2V methods to V2I setups, while addressing communication speed and volume challenges. V2X communication, which encompasses V2V, V2I, and interactions with other entities such as pedestrians, networks, and IoT devices, introduces added complexity related to trust, data synchronization, and interoperability. Future research must move beyond simulations and prioritize real-world testing to validate V2X reliability and scalability in dynamic traffic systems.
  11. 11.
    Integration of Prediction Models.
    Trajectory prediction models find integration in both fully autonomous vehicles and modern advanced driver-assistance systems (ADAS), contributing significantly to road safety. In full-stack autonomy, prediction modules inform planning about potential collisions. Automotive companies like Waymo and Cruise integrate multi-agent forecasting in their planning pipelines [^9] https://waymo.com/research/motionlm/. In ADAS [^10] https://www.mobileye.com/blog/ prediction models help infer lane-change intent and pedestrian movement, enabling early warnings that give human drivers more time to maneuver and avoid collisions. Another application involves integrating prediction with emergency braking systems, as seen in Mercedes-Benz vehicles, which use predictive cues to autonomously stop when drivers respond too late [^11] https://group.mercedes-benz.com/innovation/product-innovation/technology/pedestrian-emergency-braking-system.html. Future research should continue on integrating prediction models in traffic management systems to enable proactive safety interventions.